OpenStudy (lgbasallote):

LGBADERIVATIVE: \[\huge y = \sqrt{1+xe^{-2x}}\]

OpenStudy (lgbasallote):

\[y^2 = 1+xe^{-2x}\]

OpenStudy (lgbasallote):

\[2yy' = 1 + x(-2e^{-2x} + e{-2x})\]

OpenStudy (lgbasallote):

\[2yy' = 1 -2xe^{-2x} + e^{-2x}\]

OpenStudy (lgbasallote):

\[y' = \frac{1-2xe^{-2x} + e^{-2x}}{2\sqrt{1+xe^{-2x}}}\]

OpenStudy (lgbasallote):

was that right?

OpenStudy (helder_edwin):

no

OpenStudy (lgbasallote):

:O

OpenStudy (lgbasallote):

that 1 again huh...

OpenStudy (lgbasallote):

wait...de javu

OpenStudy (lgbasallote):

i just made the same mistake i did in http://openstudy.com/updates/4ff8d1a7e4b058f8b76342dd

OpenStudy (lgbasallote):

why can i never get that \[\frac{d}{dx} (1) = 0\] :/

OpenStudy (lgbasallote):

\[y' = \frac{-2xe^{-2x} + e^{-2x}}{2\sqrt{1+xe^{-2x}}}\] now is it right?

OpenStudy (zarkon):

yes

OpenStudy (lgbasallote):

lol i seriously hate this d/dx of 1

OpenStudy (zarkon):

I usually get quite a few of my students when I ask something like this... \[\frac{d}{dx}(\pi^2)\]

OpenStudy (zarkon):

some like to write \(2\pi\) :)

OpenStudy (helder_edwin):

when compution a derivative just remember to ignore constants that stand alone

OpenStudy (lgbasallote):

yeah i know...but everytime i do that technique i write 1 haha